Suppose the sales of a particular brand of appliances satisfy the relationship:
S(x) = 140x + 100
where S(x) represents the number of sales in year x, with x = 0 corresponding to 1982. Find the number of sales in 1987

Answer by Nate(3500)

S(5) = 140(5) + 100 = 700 + 100 = 800